find the area of the region bounded by the curves x=y³ and y=x²

1
Expert's answer
2022-05-31T11:00:59-0400

We can rewrite the equation of the first curve "x=y^3" as "\sqrt[3]{x}=y" or "y=\sqrt[3]{x}."

Let`s construct the graphical representation of the area:



The area of the region bounded by the curves equals a definite integral between the points of intersection of these curves: (0, 0) and (1, 1).


"A=\int_0^1 \sqrt[3]{x}dx - \int_0^1 x^2dx= \frac{3x\sqrt[3]{x}}{4}|_0^1 - \frac{x^3}{3}|_0^1="


"=(\frac{3}{4}-0) - (\frac{1}{3}-0)=\frac{3}{4}-\frac{1}{3}=\frac{9}{12}-\frac{4}{12}=\frac{5}{12}."


Answer: "\frac{5}{12}."